# Long distance propagation of wave beams in paraxial regime

### v1

@@ 4.2 Convergence of moments: from paraxial to Itô-Schrödinger
×$\times\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{\mathbb{R}^{2d}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}(\langle k\rangle\langle k^{\prime}\rangle)^{(p+q)\alpha}\hat{\mathfrak{C}}^{\theta}(s-s_{m},k)\hat{\mathfrak{C}}^{\theta}(s^{\prime}-s_{n},k^{\prime})\frac{\mathrm{d}k\mathrm{d}k^{\prime}}{(2\pi)^{2d}}\mathrm{d}s^{\prime}\mathrm{d}s$
×$\times\int\limits_{[0,s^{\prime}]^{2N}_{<}}\int\limits_{\mathbb{R}^{(N-1)d}}\sum\limits_{P^{2N-2}}\prod\limits_{\begin{subarray}{c}(j^{\prime},l^{\prime})\in P^{2N-2}\\ j^{\prime},l^{\prime}\neq m,n\end{subarray}}\langle k_{j^{\prime}}\rangle^{(p+q)\alpha}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
Let $\mathfrak{C}_{n,\alpha}(s)=\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{n\alpha}\hat{\mathfrak{C}}(s,k)\frac{\mathrm{d}k}{(2\pi)^{d}}$ with $\mathfrak{C}^{\theta}_{n,\alpha}(s)=\mathfrak{C}_{n,\alpha}\big{(}\frac{\eta s}{\varepsilon\theta}\big{)}$. Then we have that
@@
×$\times\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2N}_{<}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{m})\mathfrak{C}_{p+q,\alpha}^{\theta}(s^{\prime}-s_{n})\sum\limits_{P^{2N-2}}\prod\limits_{\begin{subarray}{c}(j,l)\in P^{2N-2}\\ j,l\neq m,n\end{subarray}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s_{j}-s_{l})\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
@@
×$\times\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}\mathfrak{C}^{\theta}_{p+q,\alpha}(s-s^{\prime})\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}^{N-1}\Big{(}\frac{1}{(\eta\varepsilon\theta)^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{1})\mathfrak{C}_{p+q,\alpha}^{\theta}(s^{\prime}-s_{2})\mathrm{d}s_{2}\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}.$
@@
$\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}\mathfrak{C}^{\theta}_{p+q,\alpha}(s-s^{\prime})\mathrm{d}s^{\prime}\mathrm{d}s\leq\frac{z}{\eta^{2}}\int\limits_{\mathbb{R}}\mathfrak{C}_{p+q,\alpha}(s)\mathrm{d}s=\frac{z}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|.$
@@
$\frac{1}{(\eta\varepsilon\theta)^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{1})\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{2})\mathrm{d}s_{2}\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s$
@@
$=\Big{(}\frac{1}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{\frac{\eta(s-s^{\prime})}{\varepsilon\theta}}^{\frac{\eta s}{\varepsilon\theta}}\mathfrak{C}_{p+q,\alpha}(s_{1})\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}.$
From assumption (9), $\mathfrak{C}_{p+q,\alpha}(s)\leq\lambda_{1}(d_{0})e^{-\tau_{R}|s|}$. So, we deduce that the above integral is bounded as
$\Big{(}\frac{1}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{\frac{\eta(s-s^{\prime})}{\varepsilon\theta}}^{\frac{\eta s}{\varepsilon\theta}}\mathfrak{C}_{p+q,\alpha}(s_{1})\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}$
$\leq\Big{(}\frac{\lambda_{1}(d_{0})}{\tau_{R}\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}e^{\frac{\tau_{R}\eta}{\varepsilon\theta}(s-s^{\prime})}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}\leq\lambda_{1}(d)\frac{\varepsilon\theta}{\tau_{R}^{2}\eta^{3}}\Big{(}\frac{z}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}\,.$
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|$ $\leq\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|\sum\limits_{N\geq 1}\frac{(p+q)^{2N+2}}{(N-1)!}C^{(N+1)(p+q)}\frac{z^{N}}{\eta^{2N}}\Big{(}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}^{N}$
$\leq c_{1}(p,q,z)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\,.$
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{2}(z,v)\|\leq c_{2}(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{1+\alpha}\psi^{\theta}(0)\|\,,$ (51)
@@
×$\times\Big{[}\sum\limits_{j,l}\big{(}2e^{\frac{i\eta}{\varepsilon}s^{\prime}[g(\xi_{j},k)-g(\zeta_{l},k)]}-e^{\frac{i\eta}{\varepsilon}[sg(\xi_{j},k)-s^{\prime}g(\zeta_{l},k)]}-e^{-\frac{i\eta}{\varepsilon}[sg(\zeta_{l},k)-s^{\prime}g(\xi_{j},k)]}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)$
@@
×$\times\Big{[}\sum\limits_{j,l}e^{\frac{i\eta}{\varepsilon}s^{\prime}[g(\xi_{j},k)-g(\zeta_{l},k)]}\big{(}2-e^{i\theta sg(\xi_{j},k)}-e^{-i\theta sg(\zeta_{l},k)}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)$
@@
$\leq\frac{\theta}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{\frac{\eta(z-s^{\prime})}{\varepsilon\theta}}\int\limits_{\mathbb{R}^{d}}|s|\hat{\mathfrak{C}}(s,k)\int\limits_{\mathbb{R}^{(p+q)d}}\Big{[}\sum\limits_{j,l}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\big{(}|g(\xi_{j},k)|+|g(\zeta_{l},k)|\big{)}|\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)|$
@@
$\leq C\frac{\theta}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{\frac{\eta(z-s^{\prime})}{\varepsilon\theta}}\int\limits_{\mathbb{R}^{d}}|s|\langle k\rangle^{2+2\alpha}\hat{\mathfrak{C}}(s,k)\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle\langle v_{n^{\prime}}\rangle\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(s^{\prime},v)\|\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s\mathrm{d}s^{\prime}$
$\leq Cz\frac{\theta}{\eta^{2}}\sup\limits_{0\leq s\leq z}\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle\langle v_{n^{\prime}}\rangle\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{d}}|s|\langle k\rangle^{2+2\alpha}\hat{\mathfrak{C}}(s,k)\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s\,.$
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|\leq c_{3}(p,q)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\,,$ (52)
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq\frac{1}{\eta\varepsilon\theta}\int\limits_{z}^{\infty}\int\limits_{0}^{z}\int\limits_{\mathbb{R}^{d}}\hat{\mathfrak{C}}^{\theta}(s-s^{\prime},k)\Big{[}2\sum\limits_{j,l}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)\|$
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq\frac{(p+q)^{2}}{\eta\varepsilon\theta}\sup\limits_{0\leq s\leq z}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{z}^{\infty}\int\limits_{0}^{z}\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{2\alpha}\hat{\mathfrak{C}}^{\theta}(s-s^{\prime},k)\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
Since $\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{2\alpha}\hat{\mathfrak{C}}(s,k)\frac{\mathrm{d}k}{(2\pi)^{d}}\leq Ce^{-\tau_{R}|s|}$, after a change of variables $\frac{\eta(s-s^{\prime})}{\varepsilon\theta}\to s^{\prime}$, we have
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq C\frac{(p+q)^{2}}{\eta^{2}}\sup\limits_{0\leq s\leq z}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{z}^{\infty}\int\limits_{\frac{\eta(s-z)}{\varepsilon\theta}}^{\frac{\eta s}{\varepsilon\theta}}e^{-\tau_{R}|s|}\mathrm{d}s^{\prime}\mathrm{d}s$
$\leq C\frac{(p+q)^{2}}{\tau_{R}\eta^{2}}\sup\limits_{0\leq s\leq z}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{z}^{\infty}e^{-\frac{\tau_{R}\eta s}{\varepsilon\theta}}\mathrm{d}s$
$\leq C(p+q)^{2}\frac{\varepsilon\theta}{\tau_{R}^{2}\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\,.$

### v2

@@ 4.2 Convergence of moments: from paraxial to Itô-Schrödinger
$\times\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{\mathbb{R}^{2d}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}(\langle k\rangle\langle k^{\prime}\rangle)^{(p+q)\alpha}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}^{\theta}(s-s_{m},k)||\hat{\mathfrak{C}}^{\theta}(s^{\prime}-s_{n},k^{\prime})|}\frac{\mathrm{d}k\mathrm{d}k^{\prime}}{(2\pi)^{2d}}\mathrm{d}s^{\prime}\mathrm{d}s$
$\times\int\limits_{[0,s^{\prime}]^{2N}_{<}}\int\limits_{\mathbb{R}^{(N-1)d}}\sum\limits_{P^{2N-2}}\prod\limits_{\begin{subarray}{c}(j^{\prime},l^{\prime})\in P^{2N-2}\\ j^{\prime},l^{\prime}\neq m,n\end{subarray}}\langle k_{j^{\prime}}\rangle^{(p+q)\alpha}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})|}\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
Let $\mathfrak{C}_{n,\alpha}(s)=\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{n\alpha}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}(s,k)|}\frac{\mathrm{d}k}{(2\pi)^{d}}$ with $\mathfrak{C}^{\theta}_{n,\alpha}(s)=\mathfrak{C}_{n,\alpha}\big{(}\frac{\eta s}{\varepsilon\theta}\big{)}$. Then we have that
@@
$\times\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2N}_{<}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{m})\mathfrak{C}_{p+q,\alpha}^{\theta}(s^{\prime}-s_{n})\sum\limits_{P^{2N-2}}\prod\limits_{\begin{subarray}{c}(j,l)\in P^{2N-2}\\ j,l\neq m,n\end{subarray}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s_{j}-s_{l})\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
@@
$\times\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}\mathfrak{C}^{\theta}_{p+q,\alpha}(s-s^{\prime})\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}^{N-1}\Big{(}\frac{1}{(\eta\varepsilon\theta)^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{1})\mathfrak{C}_{p+q,\alpha}^{\theta}(s^{\prime}-s_{2})\mathrm{d}s_{2}\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}.$
@@
$\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}\mathfrak{C}^{\theta}_{p+q,\alpha}(s-s^{\prime})\mathrm{d}s^{\prime}\mathrm{d}s\leq\frac{z}{\eta^{2}}\int\limits_{\mathbb{R}}\mathfrak{C}_{p+q,\alpha}(s)\mathrm{d}s=\frac{z}{\eta^{2}}{\color[rgb]{0,0,0}\int_{\mathbb{R}^{d+1}}\langle k\rangle^{(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k}.$
@@
$\frac{1}{(\eta\varepsilon\theta)^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{1})\mathfrak{C}_{p+q,\alpha}^{\theta}(s{\color[rgb]{0,0,0}{}^{\prime}}-s_{2})\mathrm{d}s_{2}\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s$
@@
$={\color[rgb]{0,0,0}\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{s_{1}}^{s}\mathfrak{C}^{\theta}_{p+q,\alpha}(s-s_{1})\mathrm{d}s^{\prime}\mathrm{d}s_{1}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\int_{\mathbb{R}^{d+1}}\langle k\rangle^{(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k\Big{)}}.$
So, we deduce that the above integral is bounded as
${\color[rgb]{0,0,0}\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{0}^{z}\int\limits_{0}^{s}(s-s_{1})\mathfrak{C}^{\theta}_{p+q,\alpha}(s_{1})\mathrm{d}s_{1}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\int_{\mathbb{R}^{d+1}}\langle k\rangle^{(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k\Big{)}}$
${\color[rgb]{0,0,0}\leq\Big{(}\frac{\varepsilon\theta z}{\eta^{3}}\int\limits_{\mathbb{R}^{d+1}}|s|\langle k\rangle^{(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k\Big{)}\Big{(}\frac{1}{\eta^{2}}\int_{\mathbb{R}^{d+1}}\langle k\rangle^{(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k\Big{)}}\,.$
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|$
$\leq\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|\sum\limits_{N\geq 1}\frac{(p+q)^{2N+2}}{(N-1)!}C^{(N+1)(p+q)}\frac{z^{N}}{\eta^{2N}}\Big{(}{\color[rgb]{0,0,0}\int\limits_{\mathbb{R}^{d+1}}\langle s\rangle\langle k\rangle^{(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k}\Big{)}^{N}$
$\leq c_{1}(p,q,z)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]}{\eta^{2}}}\,.$
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{2}(z,v)\|\leq c_{2}(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{1+\alpha}\psi^{\theta}(0)\|\,$, (51)
@@
$\times\Big{[}\sum\limits_{j,l}\big{(}2e^{\frac{i\eta}{\varepsilon}s^{\prime}[g(\xi_{j},k)-g(\zeta_{l},k)]}-e^{\frac{i\eta}{\varepsilon}[sg(\xi_{j},k)-s^{\prime}g(\zeta_{l},k)]}-e^{-\frac{i\eta}{\varepsilon}[sg(\zeta_{l},k)-s^{\prime}g(\xi_{j},k)]}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)$
@@
$\times\Big{[}\sum\limits_{j,l}e^{\frac{i\eta}{\varepsilon}s^{\prime}[g(\xi_{j},k)-g(\zeta_{l},k)]}\big{(}2-e^{i\theta sg(\xi_{j},k)}-e^{-i\theta sg(\zeta_{l},k)}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)$
@@
$\leq\frac{\theta}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{\frac{\eta(z-s^{\prime})}{\varepsilon\theta}}\int\limits_{\mathbb{R}^{d}}|s|{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}(s,k)|}\int\limits_{\mathbb{R}^{(p+q)d}}\Big{[}\sum\limits_{j,l}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\big{(}|g(\xi_{j},k)|+|g(\zeta_{l},k)|\big{)}|\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)|$
@@
$\leq C\frac{\theta}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{\frac{\eta(z-s^{\prime})}{\varepsilon\theta}}\int\limits_{\mathbb{R}^{d}}|s|\langle k\rangle^{2+2\alpha}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}(s,k)|}\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle\langle v_{n^{\prime}}\rangle\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(s^{\prime},v)\|\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s\mathrm{d}s^{\prime}$
$\leq Cz\frac{\theta}{\eta^{2}}\sup\limits_{0\leq s\leq z}\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle\langle v_{n^{\prime}}\rangle\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(s,v)\|{\color[rgb]{0,0,0}\int\limits_{\mathbb{R}^{d+1}}|s|\langle k\rangle^{2+2\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k}\,.$
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|\leq c_{3}(p,q)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]}{\eta^{2}}}\,$, (52)
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq\frac{1}{\eta\varepsilon\theta}\int\limits_{z}^{\infty}\int\limits_{0}^{z}\int\limits_{\mathbb{R}^{d}}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}^{\theta}(s-s^{\prime},k)|}\Big{[}2\sum\limits_{j,l}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)\|$
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq\frac{(p+q)^{2}}{\eta\varepsilon\theta}\sup\limits_{0\leq s\leq z}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|{\color[rgb]{0,0,0}\int\limits_{0}^{z}\int\limits_{z}^{\infty}\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{2\alpha}|\hat{\mathfrak{C}}^{\theta}(s-s^{\prime},k)|\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s\mathrm{d}s^{\prime}}\,.$
After a change of variables $s-s^{\prime}\to s$, we have
$\int_{0}^{z}\int_{z}^{\infty}|\hat{C}^{\theta}(s-s^{\prime},\cdot)|\mathrm{d}s\mathrm{d}s^{\prime}$ $=\Big{(}\int_{0}^{z}\int_{z-s^{\prime}}^{z}+\int_{0}^{z}\int_{z}^{\infty}\Big{)}|\hat{C}^{\theta}(s,k)|\mathrm{d}s\mathrm{d}s^{\prime}$
$=\int_{0}^{z}s|\hat{C}^{\theta}(s,\cdot)|\mathrm{d}s+z\int_{z}^{\infty}|\hat{C}^{\theta}(s,\cdot)|\mathrm{d}s\leq\int_{0}^{\infty}s|\hat{C}^{\theta}(s,\cdot)|\mathrm{d}s\,.$
From a change of variables $\frac{\eta s}{\varepsilon\theta}\to s$, we have
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq C(p+q)^{2}\frac{\varepsilon\theta}{\eta^{3}}\sup\limits_{0\leq s\leq z}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{\mathbb{R}^{d+1}}|s|\langle k\rangle^{2\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k\,.$